Plasma Learning is an e-learning business that specialises in providing online learning resources in accounting, economics, and marketing management. The adjusted trial balance of Plasma Learning as at 30 June 2019 as follows.

Debit $

Credit $

Cash at Bank

18500

Accounts Receivable

7800

Office Supplies

900

Prepaid Rent

3000

Equipment

60000

Accumulated Depreciation - Equipment

300

Furniture

27000

Accumulated Depreciation - Furniture

450

Accounts Payable

22300

Salary Payable

1350

Interest Payable

150

Unearned Service Revenue

2400

Loan Payable

60000

Owner’s, Capital

10000

Owner’s, Drawings

1500

Service Revenue

29600

Rent Expense

2100

Salary Expense

4700

Supplies Expense

150

Depreciation Expense - Building

300

Depreciation Expense - Furniture

450

Interest Expense

150

126550

126550

Required:

a)     Prepare Plasma Learning’s Income Statement for the year ended 30 June 2019. (8 Marks)

b)    Prepare Plasma Learning’s Statement of Changes in Equity for the year ended 30 June 2019. (4 Marks)

c)     Calculate Plasma Learning’s Profit Margin and Return on Assets ratios (the total assets at 30 June 2018 was $95000). (Show all calculations, Round off the answers to the nearest 2 decimal places). (2 Marks)

d)    (i) Calculate Plasma Learning’s Current ratio and the Debt-to-total assets ratio at 30 June 2019. (Show all calculations, Round off the answers to the nearest 2 decimal places). (2 Marks)

(ii) Discuss Plasma Learning’s liquidity and solvency positions at 30 June 2019 as compared to 2018 (the Current ratio was 1.85:1 and the Debt-to-total assets ratio was 53.87% at 30 June 2018). (4 Marks) (Total 20 Marks)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Plasma Learning
Answer a
Income Statement Amount $
Service Revenue        29,600.00
Less:
Rent Expense          2,100.00
Salaries Expense          4,700.00
Supplies Expense             150.00
Depreciation Expense - Building             300.00
Depreciation Expense - Furniture             450.00
Interest Expense             150.00
Net profit       21,750.00
Answer b
Statement of changes in equity Amount $
Opening Balance        10,000.00
Add: Net Income        21,750.00
Less: Drawings          1,500.00
Closing Balance       30,250.00
Balance Sheet Amount $ Amount $
Current Assets
Cash        18,500.00
Accounts Receivable          7,800.00
Office Supplies             900.00
Prepaid rent          3,000.00
Total Current Assets       30,200.00
Non Current Assets
Equipment        60,000.00
Accumulated Depreciation - Equipment            (300.00)        59,700.00
Furniture        27,000.00
Accumulated Depreciation - Furniture            (450.00)        26,550.00
Fixed Assets       86,250.00
Total Assets     116,450.00
Liabilities & Owner's Equity
Liabilities Amount $ Amount $
Current Liabilities
Accounts Payable        22,300.00
Salary Payable          1,350.00
Interest Payable             150.00
Unearned Service Revenue          2,400.00
Current Liabilities       26,200.00
Non- Current Liabilities
Loan Payable        60,000.00
Non- Current Liabilities       60,000.00
Total Liabilities       86,200.00
Owner's Equity        30,250.00
Owner's Equity       30,250.00
Total Liabilities & Owner's Equity     116,450.00
Answer c Note
Net profit        21,750.00 A
Service Revenue        29,600.00 B
Profit Margin 73.48% C=A/B
Opening total assets        95,000.00 D
Closing total assets     116,450.00 E
Average total assets     105,725.00 F=(D+E)/2
Net profit        21,750.00 See A
Return on Assets 20.57% G=A/F
Answer d i
Total Current Assets        30,200.00 H
Current Liabilities        26,200.00 I
Current ratio                  1.15 J=H/I
Answer d i
Current ratio of previous year was 1.85:1 which is greater than currant year so the liquidity of the firm has worsened.
Answer d i
Total Liabilities        86,200.00 K
Total Assets     116,450.00 See E
Debt-to-total assets ratio 74.02% L=K/E
Answer d ii
Debt-to-total assets ratio previous year was 53.87% which is lesser than currant year so the solvency of the firm has worsened.